OpenStudy (anonymous):

Simplify log5 (x+2) - log5 10= log 5 100 a.108 b.8 c.998 d.1000

OpenStudy (asnaseer):

I think you mean solve for x rather than simplify?

OpenStudy (asnaseer):

do you know the various log rules?

OpenStudy (anonymous):

Yes, i do but how do i apply it to this?

OpenStudy (asnaseer):

start with the left-hand-side

OpenStudy (asnaseer):

use the rule:\[\log_a(p)-\log_a(q)=\log_a(\frac{p}{q})\]

OpenStudy (anonymous):

but how does this formula pertain to the question?

OpenStudy (asnaseer):

its asking you to find the value of x in this equation. so you need to first simplify the left-hand-side

OpenStudy (asnaseer):

you have this on the left-hand-side:\[\log_5(x+2) - \log_5(10)\]use the rule I gave above to first simplify this expression

OpenStudy (anonymous):

thank you so muc, i have a question in order to simplify it what do i do with the log5 10 ?

OpenStudy (asnaseer):

use the rule I gave above, i.e use:\[\log_a(p)-\log_a(q)=\log_a(\frac{p}{q})\]in your case p=x+2 and q=10

OpenStudy (asnaseer):

and 'a' is the base of the log which is 5 in your case

OpenStudy (anonymous):

Ohhhhh, the answer i have come up with is 108 Is that correct?

OpenStudy (asnaseer):

no

OpenStudy (asnaseer):

tell me what you got when you simplified the left-hand-side please

OpenStudy (anonymous):

I give up, Thank You so much for your help though!

OpenStudy (asnaseer):

I take you through it step-by-step if you want to learn?

OpenStudy (asnaseer):

*I can ...

OpenStudy (anonymous):

okay sure then

OpenStudy (asnaseer):

ok, so you are given:\[log_5(x+2) - \log_5(10)= \log_5(100)\]use the rule I gave above, we can simplify the left-hand-side as follows:\[\log_5(\frac{x+2}{10})=\log_5(100)\]

OpenStudy (asnaseer):

do you follow so far?

OpenStudy (anonymous):

so i jsut plug the numbers in then?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

no - no need to plug in any numbers

OpenStudy (asnaseer):

next step is to notice that both sides are logs of something to base 5

OpenStudy (anonymous):

so i cancel them out? or add them up?

OpenStudy (asnaseer):

therefore the "something" within the logs must be equal

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

therefore:\[\frac{x+2}{10}=100\]

OpenStudy (asnaseer):

do you know how to find x from here?

OpenStudy (anonymous):

obviously, lol okay simplify

OpenStudy (anonymous):

YES THANK YOU SO MUCH!

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

you are a mathj genius!

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

the answer is c, 998 right.

OpenStudy (asnaseer):

correct - well done! :)

OpenStudy (anonymous):

*claps* Thank You.

OpenStudy (asnaseer):

:D